Solve the system
2x + 3y = 4
x + 4y = 7

Answer by edjones(8007)

2x + 3y = 4
x + 4y = 7
subtract 4y from both sides of the 2nd equation. x=7-4y
rplace x with 7-4y in the 1st equation. 2(7-4y)+3y=4
14-8y+3y=4
14-5y=4
subtract 14 from both sides. -5y=-10
divide -5 into both sides. y=2
replace y with 2 in either equation. x+8=7
subtract 8 from both sides. x=-1
Ans. x=-1, y=2
Check:
2*-1+3*2=4 ...-2+6=4
-1+8=7
Ed